# Deflection angle in the strong deflection limit: a perspective from local geometrical invariants and matter distributions

### v1

@@ III Derivation of the deflection angle in the strong deflection limit
where we have used $\lim_{r_{0}\to r_{\mathrm{m}}}(V^{\prime}_{0}/R^{\prime}_{0})=V_{\mathrm{m}}^{\prime\prime}/R^{\prime\prime}_{\mathrm{m}}$. By inverting Eq. (28), we express the above expansions in terms of $(b/b_{\mathrm{c}}-1)$ as follows
@@
These results lead to the deflection angle (34) with the following expression for the strong field limit coefficients:
@@
Thus, when $R^{\prime}_{\mathrm{m}}=0$ (i.e., the photon sphere coincides with the wormhole throat), the strong field limit coefficients arising from the divergence part can be expressed solely in terms of local quantities. In particular, $\bar{a}$ remains entirely determined by $V^{\prime\prime}_{\mathrm{m}}$ and retains the same form as in the case $R^{\prime}_{\mathrm{m}}\neq 0$, while $\bar{b}$ is given by
@@ IV Formula for the deflection angle in the strong deflection limit
To facilitate a definite physical interpretation of the subsequent analysis, we introduce the following orthonormal tetrad:
@@
In this tetrad basis, the components of the Einstein tensor $G_{(\mu)(\nu)}=G_{ab}e_{(\mu)}^{a}e_{(\nu)}^{b}$ are given by
@@
with all other components vanishing identically, where the prime denotes differentiation with respect to the radial coordinate $r$. These tetrad components of the Einstein tensor encode the curvature as measured by static observers, thereby providing a coordinate-invariant characterization of the local geometry. The contracted Bianchi identity, $\nabla_{a}G^{ab}=0$, in the orthonormal frame reads
@@
Furthermore, by using the Einstein equations, the derivatives of the metric functions can be expressed as
@@
These relations provide a convenient formulation for the metric functions in terms of the Einstein tensor components.
First, we focus on $G_{(1)(1)}$ at the photon sphere (if it exists), which is located at $r=r_{\mathrm{m}}$. According to Eqs. (5) and (13), the value of $G_{(1)(1)}$ at the photon sphere is given by
$G_{(1)(1)}^{\mathrm{m}}=\frac{2\left[\>\!R_{\mathrm{m}}-3m(r_{\mathrm{m}})\>\!\right]}{R_{\mathrm{m}}^{3}}.$ (55)
Note that this relation establishes a direct link between the local geometrical quantity $G_{(1)(1)}^{\mathrm{m}}$ and the gravitational mass $m(r_{\mathrm{m}})$, which represents the total mass contained within a sphere of radius $r_{\mathrm{m}}$. The sign of $G_{(1)(1)}^{\mathrm{m}}$ is positive when $R_{\mathrm{m}}>3m(r_{\mathrm{m}})$, negative when $R_{\mathrm{m}}<3m(r_{\mathrm{m}})$, and vanishes when $R_{\mathrm{m}}=3m(r_{\mathrm{m}})$. Moreover, the value $G_{(1)(1)}^{\mathrm{m}}$ is independent of the dynamical stability of photon circular orbits and thus remains valid even for anti-photon spheres [30, 31, 32].
In a similar manner, we can relate $G_{(0)(0)}$ at the the photon sphere to the radial variation of the Misner–Sharp mass. Using Eq. (48), we have
@@
Now, we establish the relation between the second derivative of the effective potential, $V^{\prime\prime}_{\mathrm{m}}$, and the tetrad components of the Einstein tensor, $G_{(\mu)(\mu)}^{\mathrm{m}}$, both evaluated at the photon sphere. Substituting Eqs. (13), (53), and (54) into Eq. (15) yields
@@
This expression demonstrates that the local spacetime geometry directly contributes to the curvature of the effective potential at the photon sphere. Furthermore, this contribution arises exclusively through the dimensionless product of the square of the areal radius $R_{\mathrm{m}}$ and the curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$, thereby underscoring the fundamental importance of this quantity. In particular, a necessary condition for unstable photon circular orbits (i.e., for $V_{\mathrm{m}}^{\prime\prime}<0$) is given by
@@
Notably, when $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$ vanishes—as it does in vacuum in general relativity—the above inequality is trivially satisfied, yielding $V^{\prime\prime}_{\mathrm{m}}=-2$. This result confirms that an anti-photon sphere cannot exist; only a photon sphere is possible.
@@
and the constant shift correction $\bar{b}$ given by
@@
Notably, these new expressions reveal that both the logarithmic divergence rate $\bar{a}$ and the constant shift correction $\bar{b}$ arising from the divergence part are solely determined by local geometrical quantities, namely, the areal radius $R_{\mathrm{m}}$ and the tetrad components $G_{(\mu)(\mu)}^{\mathrm{m}}$. Specifically, $\bar{a}$ depends only on $R_{\mathrm{m}}$ and the curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$, which defines the scaling via the dimensionless product $R_{\mathrm{m}}^{2}(G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$. Consequently, this combination essentially governs the rate of the logarithmic divergence of the strong deflection angle in the strong deflection limit.
On the other hand, the first term in $\bar{b}$—the constant contribution of the divergent integral—is determined solely by the dimensionless products $R_{\mathrm{m}}^{2}(G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$ (through $\bar{a}$) and, depending on the case, by either $R_{\mathrm{m}}^{2}G_{(1)(1)}^{\mathrm{m}}$ for $R^{\prime}_{\mathrm{m}}\neq 0$ or $R_{\mathrm{m}}^{2}G_{(0)(0)}^{\mathrm{m}}$ for $R^{\prime}_{\mathrm{m}}=0$. Although, as discussed in the previous section, the separation between the divergent and regular parts depends on the choice of the expansion parameter $z$, the present separation—with its first term expressed purely in terms of local quantities—remains highly valuable and proves effective for estimating these local quantities, as will be discussed immediately below. The remaining contribution to $\bar{b}$ arises from the regular part $I_{\mathrm{R}}(r_{\mathrm{m}})$, which is evaluated within a specific spacetime model.
Consequently, these results imply that by measuring the deflection angle in the strong deflection regime, we can extract valuable information about the curvature in the vicinity of the photon sphere. If $\bar{a}$ and $R_{\mathrm{m}}$ are determined from observations, the specific curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$ can be directly derived. Moreover, if $\bar{b}$ is also measured and a specific spacetime model is assumed—so that all the relevant quantities can be computed theoretically for direct comparison with observations—then $G_{(1)(1)}^{\mathrm{m}}$ can be deduced for the case $R^{\prime}_{\mathrm{m}}\neq 0$ and $m(r_{\mathrm{m}})$ determined via Eq. (55), or then $G_{(0)(0)}^{\mathrm{m}}$ can be deduced for the case $R^{\prime}_{\mathrm{m}}=0$ and $\mathrm{d}m/\mathrm{d}R|_{r=r_{\mathrm{m}}}$ determined via Eq. (56). In other words, if the theoretical predictions for $\bar{a}$ and $\bar{b}$ agree reasonably well with the observed values, this approach enables us to infer the local curvature and the gravitational mass, thereby providing a robust consistency check of the underlying gravitational model.
@@ V Interpreting strong deflection in terms of local matter fields
In the previous section, we derived the logarithmic divergence of the deflection angle in the strong deflection limit and reexpressed both its divergence rate and constant shift contribution in terms of local curvature quantities, specifically the tetrad components of the Einstein tensor. Importantly, our formulation is independent of any specific theoretical framework, making it applicable not only to general relativity but also to alternative theories of gravity.
In this section, we aim to elucidate the relationship between the deflection angle in the strong deflection limit and the matter field quantities. According to the Einstein equations, the matter fields are related to the spacetime curvatures as follows:
@@
where $\rho(r)$ denotes the local energy density, $P(r)$ the radial pressure, and $\Pi(r)$ the tangential pressure as measured by static observers. Note that while we use the standard terminology for matter fields in general relativity, these expressions can be recast in terms of the corresponding curvature quantities via Eq. (62), thus extending the applicability of our discussion to other gravitational theories.
@@
$\rho_{\mathrm{m}}=\rho(r_{\mathrm{m}}),\quad P_{\mathrm{m}}=P(r_{\mathrm{m}}),\quad\Pi_{\mathrm{m}}=\Pi(r_{\mathrm{m}}).$ (63)
@@
Finally, we express the strong field limit coefficients $\bar{a}$ and $\bar{b}$ in terms of these matter field quantities. Substituting Eqs. (62) into Eqs. (60) and (68), we obtain
@@
These expressions demonstrate that $\bar{a}$ depends solely on $R_{\mathrm{m}}^{2}\big{(}\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\big{)}$, while the first term in $\bar{b}$ is determined by both $R_{\mathrm{m}}^{2}\big{(}\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\big{)}$ (through $\bar{a}$) and, depending on the case, either $R_{\mathrm{m}}^{2}P_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}\neq 0$) or $R_{\mathrm{m}}^{2}\rho_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}=0$). Thus, we conclude that $\bar{a}$ and $\bar{b}$ are determined exclusively by the local geometric quantity $R_{\mathrm{m}}$ and by specific combinations of matter field variables, namely, $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}$ and either $P_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}\neq 0$) or $\rho_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}=0$). Based on the discussion at the end of Sec. III, measurements of the deflection angle in the strong deflection regime allow us to infer the combination $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}$ and either the radial pressure $P_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}\neq 0$) or the energy density $\rho_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}=0$), thereby directly linking observable lensing phenomena to the local matter distribution at the photon sphere.
In the following, we examine the relationship between the properties of the matter field quantities defined above and the strong field limit coefficients. Detailed calculations of the strong field limit coefficients for specific spacetime models have been extensively addressed in the literature, and we refer the reader to those works for further details. Nevertheless, it is instructive to present results for several well-known spacetimes. Furthermore, we aim to elucidate the universal properties that emerge when particular combinations of the matter field quantities vanish.
As the simplest case, we consider a vacuum spacetime in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Schwarzschild solution
@@
where $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\>\!\mathrm{d}\varphi^{2}$. Here, $M$ denotes the constant mass enclosed within the vacuum region, and the standard gauge $R=r$ is adopted. In the vacuum region, all matter field quantities vanish; hence, at $r_{\mathrm{m}}=3M$, we obtain
@@
with the critical impact parameter given by $b_{\mathrm{c}}=3\sqrt{3}M$. Note that $\bar{a}=1$ is a direct consequence of the vanishing matter fields in a vacuum spacetime. These results recover the well-known expressions for the Schwarzschild solution [5, 6] because our definition of $z$ is equivalent to that used in the references in the vacuum case. This confirms that our formulation is consistent with the standard vacuum case in general relativity.
Next, we consider an electrovac spacetime (i.e., a vacuum spacetime containing only electromagnetic fields) in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Reissner–Nordström solution
@@
where $M$ and $Q$ denote the constant mass and electric charge characterizing the electrovac region, respectively, and the standard gauge $R=r$ is adopted. In the electrovac, matter field quantities are given by
@@
where $r_{\mathrm{m}}=(3M+\sqrt{9M^{2}-8Q^{2}})/2$ and $b_{\mathrm{c}}^{2}=3r_{\mathrm{m}}^{2}/(1-Q^{2}/r_{\mathrm{m}}^{2})$. Note that these results agree with those in Refs. [9, 10], which employ the same expansion parameter as used here. These expressions clearly indicate that both coefficients depend solely on the dimensionless parameter $Q/r_{\mathrm{m}}$.
Next, consider a special case in which the spacetime is nonvacuum but satisfies $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$. In this situation, the strong field limit coefficient $\bar{a}$ remains equal to $1$, thereby mirroring the Schwarzschild result despite the presence of nonvanishing matter fields. This demonstrates that when the matter fields satisfy $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$, the strong deflection limit exhibits universal behavior with $\bar{a}=1$, irrespective of the specific details of the matter content.
@@
This explicitly shows that $\rho+\Pi=0$ throughout the spacetime, which naturally leads to a universal behavior in the strong deflection limit, ensuring that the strong field limit coefficient remains $\bar{a}=1$. In other words, even though the spacetime is not strictly vacuum due to the presence of the massless scalar field, the specific structure of its energy-momentum tensor guarantees that $\bar{a}=1$, thereby retaining the Schwarzschild value.
@@
Notably, this matter configuration satisfies $\rho+\Pi=0$ throughout the spacetime, thereby leading to the universal behavior observed in the strong deflection limit. Indeed, the strong field limit coefficients for the JNW spacetime are given by
@@
where $b_{\mathrm{c}}=(r_{\mathrm{g}}/2)(2\gamma+1)^{(2\gamma+1)/2}(2\gamma-1)^{-(2\gamma-1)/2}$. Thus, even in the presence of a nontrivial scalar field, if the matter satisfies $\rho+\Pi=0$, the strong deflection limit exhibits the universal behavior characterized by $\bar{a}=1$. It should be noted that the first term in $\bar{b}$ differs from that in Refs. [5, 12] due to a different choice of the parameter $z$.
@@
where $a$ is a constant parameter. This solution describes a traversable wormhole with its throat located at $r=0$ (i.e., $R=a$), supported by a phantom massless scalar field, and it possesses the photon sphere at $r=0$. Therefore, $R_{\mathrm{m}}^{\prime}=0$ holds in this case. The associated scalar field quantities are given by
@@
which agree with those in Ref. [14], employing the same expansion parameter as used here. Note that for the Ellis wormhole under consideration, the conditions $R^{\prime\prime\prime}_{\mathrm{m}}=0$ and $V^{\prime\prime\prime}_{\mathrm{m}}=0$ uniquely satisfied, which causes the coefficient of $(b/b_{\mathrm{c}}-1)^{1/2}$ in $c_{2}$ to vanish. Consequently, the error term in Eq. (82) is of order $(b/b_{\mathrm{c}}-1)\log(b/b_{\mathrm{c}}-1)$.

### v2

@@ III.2 $R^{\prime}_{\mathrm{m}}=0$
where we have used $\lim_{r_{0}\to r_{\mathrm{m}}}(V^{\prime}_{0}/R^{\prime}_{0})=V_{\mathrm{m}}^{\prime\prime}/R^{\prime\prime}_{\mathrm{m}}$. By inverting Eq. (28), these expansions can be expressed in terms of $(b/b_{\mathrm{c}}-1)$ as follows:
@@
These results lead to the deflection angle in Eq. (34) with the following expression for the strong field limit coefficients:
@@
Thus, when $R^{\prime}_{\mathrm{m}}=0$ (i.e., when the photon sphere coincides with the wormhole throat), the strong field limit coefficients arising from $I_{\mathrm{D}}$ can be expressed solely in terms of local quantities. In particular, $\bar{a}$ remains entirely determined by $V^{\prime\prime}_{\mathrm{m}}$ and retains the same form as in the $R^{\prime}_{\mathrm{m}}\neq 0$ case, while $\bar{b}$ is given by
@@ IV Formula for the deflection angle in the strong deflection limit
To facilitate a definite physical interpretation of the subsequent analysis, we introduce the following orthonormal basis:
@@
Then, the tetrad components of the Einstein tensor $G_{(\mu)(\nu)}=G_{ab}e_{(\mu)}^{a}e_{(\nu)}^{b}$ are given by
@@
with all other components vanishing identically, where the prime denotes differentiation with respect to the radial coordinate $r$. These quantities $G_{(\mu)(\nu)}$ encode curvature as measured by static observers, thereby providing a coordinate-invariant characterization of the local geometry. The contracted Bianchi identity, $\nabla_{a}G^{ab}=0$, in the tetrad frame reads
@@
Furthermore, from Eqs. (48) and (49), the derivatives of the metric functions can be expressed as
@@
Now, we focus on $G_{(1)(1)}$ at $r=r_{\mathrm{m}}$ (assuming that the photon sphere exists). According to Eqs. (5) and (13), this value is given by
$G_{(1)(1)}^{\mathrm{m}}=\frac{2\left[\>\!R_{\mathrm{m}}-3m(r_{\mathrm{m}})\>\!\right]}{R_{\mathrm{m}}^{3}},$ (55)
which establishes a direct link between the local geometrical quantity $G_{(1)(1)}^{\mathrm{m}}$ and the gravitational mass $m(r_{\mathrm{m}})$. Here, the value $G_{(1)(1)}^{\mathrm{m}}$ is positive when $R_{\mathrm{m}}>3m(r_{\mathrm{m}})$, negative when $R_{\mathrm{m}}<3m(r_{\mathrm{m}})$, and vanishes when $R_{\mathrm{m}}=3m(r_{\mathrm{m}})$. Moreover, $G_{(1)(1)}^{\mathrm{m}}$ is independent of the dynamical stability of photon circular orbits and remains valid even for anti-photon spheres [33, 34, 35].
Similarly, we can relate $G_{(0)(0)}$ at $r=r_{\mathrm{m}}$ to the radial variation of the Misner–Sharp mass. Using Eq. (48), we have
@@
Now, we establish the relation between $V^{\prime\prime}_{\mathrm{m}}$ and $G_{(\mu)(\mu)}^{\mathrm{m}}$, both evaluated at the photon sphere. Substituting Eqs. (13), (52), and (54) into Eq. (15) yields
@@
This expression demonstrates that the local spacetime geometry directly contributes to $V^{\prime\prime}_{\mathrm{m}}$. Notably, this contribution arises exclusively through the dimensionless product $R_{\mathrm{m}}^{2}(G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$, thereby underscoring its fundamental importance. In particular, a necessary condition for the existence of unstable photon circular orbits (i.e., $V_{\mathrm{m}}^{\prime\prime}<0$) is
@@
and the constant offset correction $\bar{b}$ given by
@@
Notably, these expressions reveal that both the logarithmic divergence rate $\bar{a}$ and the constant offset correction $\bar{b}$ arising from $I_{\mathrm{D}}$ are determined solely by local geometrical quantities, i.e., $R_{\mathrm{m}}$ and $G_{(\mu)(\mu)}^{\mathrm{m}}$. Specifically, $\bar{a}$ depends only on $R_{\mathrm{m}}$ and the curvature sum $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$, which enters as the dimensionless product $R_{\mathrm{m}}^{2}(G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$. Consequently, this combination essentially governs the rate of the logarithmic divergence of the strong deflection angle.
The decomposition of $\bar{b}$ shows that its first term is determined solely by the dimensionless products $R_{\mathrm{m}}^{2}(G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$ (through $\bar{a}$) and, depending on the case, either by $R_{\mathrm{m}}^{2}G_{(1)(1)}^{\mathrm{m}}$ for $R^{\prime}_{\mathrm{m}}\neq 0$ or by $R_{\mathrm{m}}^{2}G_{(0)(0)}^{\mathrm{m}}$ for $R^{\prime}_{\mathrm{m}}=0$. Although, as discussed in the previous section, the decomposition between the divergent and regular parts depends on the choice of the expansion parameter $z$, the particular decomposition—with its first term expressed purely in terms of local quantities—remains highly valuable for estimating these local quantities. The remaining contribution to $\bar{b}$ arises from the regular part $I_{\mathrm{R}}(r_{\mathrm{m}})$, which is evaluated within a specific spacetime model.
Consequently, these results imply that we can extract information about the curvature near the photon sphere by measuring the deflection angle in the strong deflection regime. Specifically, if $\bar{a}$ and $R_{\mathrm{m}}$ are determined from observations, the curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$ can be directly derived. Moreover, if $\bar{b}$ is also measured and a specific spacetime model is assumed—so that all the relevant quantities can be computed for direct comparison with observations—then we can deduce $G_{(1)(1)}^{\mathrm{m}}$ (with $m(r_{\mathrm{m}})$) determined via Eq. (55) in the case $R^{\prime}_{\mathrm{m}}\neq 0$, or $G_{(0)(0)}^{\mathrm{m}}$ (with $\mathrm{d}m/\mathrm{d}R|_{r=r_{\mathrm{m}}}$) determined via Eq. (56) in the case $R^{\prime}_{\mathrm{m}}=0$. In other words, if the theoretical predictions for $\bar{a}$ and $\bar{b}$ agree reasonably well with the observed values, this approach enables us to infer the local curvature and the gravitational mass, thereby providing a robust consistency check of the underlying gravitational model.
@@ V Interpreting strong deflection in terms of local matter fields
In this section, we elucidate the relationship between the deflection angle in the strong deflection limit and matter field quantities. To this end, we adopt the following definitions:
@@
where, in general relativity, $\rho(r)$ denotes the local energy density, $P(r)$ the radial pressure, and $\Pi(r)$ the tangential pressure as measured by static observers. In alternative gravitational theories, however, $\rho$, $P$, and $\Pi$ may be interpreted as components of a more general tensor encoding both matter distribution and additional curvature contributions. Note that, while we use the standard terminology for matter fields in general relativity, these expressions can be recast in terms of the corresponding curvature quantities via Eq. (62), thus extending the applicability of our discussion to other gravitational theories.
@@
$\rho_{\mathrm{m}}\equiv\rho(r_{\mathrm{m}}),\quad P_{\mathrm{m}}\equiv P(r_{\mathrm{m}}),\quad\Pi_{\mathrm{m}}\equiv\Pi(r_{\mathrm{m}}).$ (63)
@@
We find that the inequality (58) for the existence of the photon sphere (i.e., $V_{\mathrm{m}}^{\prime\prime}<0$) can be rewritten as
$8\pi R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\right)<1,$ (67)
which implies that the sum $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}$ must be bounded from above for a photon sphere to exist. Notably, when $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}$ vanishes—as it does, for example, in vacuum in general relativity—the inequality is trivially satisfied, yielding $V^{\prime\prime}_{\mathrm{m}}=-2$. In this case, the result shows that a photon sphere can exist, whereas an anti-photon sphere cannot.
Finally, we express the strong field limit coefficients $\bar{a}$ and $\bar{b}$ in terms of these matter field quantities. By substituting Eqs. (62) into Eqs. (60) and (69), we obtain
@@
These expressions demonstrate that the divergent part contributions to the strong field limit coefficients are entirely determined by the local geometric scale $R_{\mathrm{m}}$ and by specific combinations of matter field variables. Specifically, $\bar{a}$ depends solely on $R_{\mathrm{m}}^{2}\big{(}\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\big{)}$, while the first term in $\bar{b}$ is further influenced, via $\bar{a}$, by either $R_{\mathrm{m}}^{2}P_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}\neq 0$) or $R_{\mathrm{m}}^{2}\rho_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}=0$). Based on the discussion at the end of Sec. III, measurements of the deflection angle in the strong deflection regime allow us to infer the combination $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}$ and either the radial pressure $P_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}\neq 0$) or the energy density $\rho_{\mathrm{m}}$ (for $R^{\prime}_{\mathrm{m}}=0$), thereby directly linking observable lensing phenomena to the local matter distribution at the photon sphere.
In the following, we examine the relationship between the properties of the matter field quantities defined above and the strong field limit coefficients. Detailed calculations of the strong field limit coefficients for specific spacetime models have been extensively addressed in the literature, and we refer the reader to those works for further details. Nevertheless, it is instructive to present results for several well-known spacetimes. Furthermore, we elucidate the universal properties that emerge when particular combinations of the matter field quantities vanish.
As the simplest case, we consider a vacuum spacetime in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Schwarzschild metric
@@
where $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\>\!\mathrm{d}\varphi^{2}$. Here, $M$ denotes the constant mass enclosed within the vacuum region, and the standard gauge $R=r$ is adopted. Since all matter field quantities vanish, at $r_{\mathrm{m}}=3M$ we obtain
@@
with the critical impact parameter given by $b_{\mathrm{c}}=3\sqrt{3}M$. Note that $\bar{a}=1$ is a direct consequence of the vanishing matter fields in a vacuum spacetime. These results recover the well-known expressions for the Schwarzschild solution [5, 6] because our definition of $z$ is equivalent to that used in those references for the vacuum case. This confirms that our formulation is fully consistent with the standard vacuum case in general relativity.
Next, we consider an electrovacuum spacetime (i.e., a vacuum spacetime containing only electromagnetic fields) in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Reissner–Nordström metric
@@
where $M$ and $Q$ denote the constant mass and electric charge characterizing the electrovacuum region, respectively, and the standard gauge $R=r$ is adopted. In the electrovacuum, matter field quantities are given by
@@
where $r_{\mathrm{m}}=(3M+\sqrt{9M^{2}-8Q^{2}})/2$ and $b_{\mathrm{c}}^{2}=3r_{\mathrm{m}}^{2}/(1-Q^{2}/r_{\mathrm{m}}^{2})$. Note that these results agree with those in Refs. [9, 10], which employ the same expansion parameter as used here. These expressions indicate that both coefficients depend solely on the dimensionless parameter $Q/r_{\mathrm{m}}$.
Next, consider a special case in which the spacetime is nonvacuum, but matter fields satisfy $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$. In this situation, the strong field limit coefficient $\bar{a}$ remains equal to $1$, thereby mirroring the Schwarzschild result despite the presence of nonvanishing matter fields. This demonstrates that when $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$, the strong deflection limit exhibits universal behavior with $\bar{a}=1$, irrespective of the specific details of the matter content.
@@
This result explicitly shows that $\rho+\Pi=0$ throughout the spacetime, which naturally leads to a universal behavior in the strong deflection limit by ensuring that $\bar{a}=1$. In other words, even though the spacetime is not strictly vacuum due to the presence of the massless scalar field, the specific structure of its energy-momentum tensor guarantees that $\bar{a}=1$, thereby retaining the Schwarzschild value.
@@
Since this matter configuration satisfies $\rho+\Pi=0$ throughout the spacetime, the strong deflection limit coefficients are given by
@@
where $b_{\mathrm{c}}=(r_{\mathrm{g}}/2)(2\gamma+1)^{(2\gamma+1)/2}(2\gamma-1)^{-(2\gamma-1)/2}$. Thus, even in the presence of a nontrivial scalar field, if $\rho+\Pi=0$, the strong deflection limit exhibits the universal behavior characterized by $\bar{a}=1$. For completeness, note that the first term in $\bar{b}$ differs from that in Refs. [5, 12] due to a different choice of the parameter $z$.
@@
where $a$ is a constant parameter. This solution describes a traversable wormhole with its throat located at $r=0$ (i.e., the areal radius $R=a$), supported by a phantom massless scalar field, and it possesses the photon sphere at $r=0$. Therefore, $R_{\mathrm{m}}^{\prime}=0$ holds in this case. The associated scalar field quantities are given by
@@
which agree with those in Ref. [14], employing the same expansion parameter used here. Note that for the Ellis wormhole under consideration, the conditions $R^{\prime\prime\prime}_{\mathrm{m}}=0$ and $V^{\prime\prime\prime}_{\mathrm{m}}=0$ are satisfied, which causes the coefficient of $(b/b_{\mathrm{c}}-1)^{1/2}$ in $c_{2}$ to vanish. Consequently, the error term in Eq. (83) is of order $(b/b_{\mathrm{c}}-1)\log(b/b_{\mathrm{c}}-1)$.